# FAST: Feature Aware Similarity Thresholding for Weak Unlearning in Black-Box Generative Models

### v1

@@ III-A Machine Unlearning
(Exact Weak Unlearning): Given a retrained generative model $G_{\theta^{r}}(.)$, we say the model $G_{\theta^{u}}(.)$ is an exact weak unlearned model iff $\forall z\sim P_{\mathcal{Z}},\mathcal{O}\subset\mathcal{X}$ the following holds:
@@
(($\epsilon,\delta$) Approximate Weak Unlearning): Given a retrained generative model $G_{\theta^{r}}(.)$, we say the model $G_{\theta^{u}}(.)$ is an $(\epsilon,\delta)$ approximate weak unlearned model for a given $\epsilon,\delta\geq 0$ iff $\forall z\sim P_{\mathcal{Z}},\mathcal{O}\subset\mathcal{X}$ the following conditions hold:
@@
III-B Filtering as Weak Unlearning
@@
Figure 1: Filtering as Weak Unlearning in Black-Box Generative Models: The left-most block represents the black-box generator with post-hoc blocking layer with parameters $\theta^{b}=(\theta^{init},t)$ whose output distribution $P_{\theta^{b}}$, the middle block represents a retrained generative model with output distribution $P_{\theta^{r}}$, the right-most block represents a $(\epsilon,0)$ unlearned model with output distribution $P_{\theta^{u}}$.
@@
Remarks:
Here the $d_{ln-TV}(.||.)$ denotes the total variational distance in terms of the log of the distributions.
Proof:
@@
Now, using ineq. (ii) and ineq. (iii) we get the below inequalities repectively
@@
IV Methodology
IV-A Problem Formulation and Method Overview
@@
As previously mentioned, the pre-trained generator $G_{\theta^{init}}$ is trained using a dataset $D$ comprising $m$ independent and identically distributed (i.i.d) samples denoted as $\{x_{i}\}_{i=1}^{m}$, where $x_{i}\overset{iid}{\sim}P_{\mathcal{X}}$. In our specific scenario, both the dataset $D$ and the initial model parameters $\theta^{init}$ remain undisclosed. Our objective is to formulate a filtering mechanism that relies on user feedback to selectively exclude undesired outputs. In this context, the user is presented with a collection of $s$ samples, denoted as $\mathcal{S}=\{y_{i}\}_{i=1}^{s}$, where each $y_{i}$ represents a generated sample originating from the pre-trained GAN. Within this user feedback system, the user’s role is to designate a relatively small subset of these samples, denoted as $\mathcal{S}_{n}=\{y^{n}_{i}\}_{i=1}^{r}$, as negative samples. These negative samples correspond to the instances that exhibit undesired features. Additionally, the user selects a portion of the samples, referred to as $\mathcal{S}_{p}=\{y^{p}_{i}\}_{i=1}^{r}$, as positive samples. Positive samples are those that do not possess the undesired features and are considered desirable. It’s essential to note that gathering user feedback can be resource-intensive. Therefore, we operate within a low-data regime, where the no of feedback samples marked by the user, denoted as $s$, is very small (constrained to be less than 100) i.e. $r\ll s$. This limitation acknowledges the potential challenges associated with obtaining extensive user feedback and emphasizes the need to devise effective filtering mechanisms even in scenarios with limited data availability.
@@
In this study, we implement a two-step process to filter out undesired outputs effectively. In the initial step, we encode the representation of the undesired feature within the latent space. This latent space may either be the implicit latent space inherent to the Generative Adversarial Network (GAN) or a latent space that is learnable. Subsequently, in the second step, we compute the projection similarity between newly sampled latent vectors and this encoded representation of the undesired feature. This similarity calculation serves as the basis for determining the degree of alignment between the latent vectors and the undesired feature. Utilizing this similarity metric, we judiciously apply a thresholding mechanism to separate and exclude outputs that contain the undesired feature.
IV-B Filtering Mechanism
IV-B1 Step-1 Undesired Feature Representation
@@
In this particular phase of our methodology, the central objective is to uncover a comprehensive and distinct representation of the undesired feature. This representation should exhibit a clear presence in the negative images while being conspicuously absent in the positive images. This task is inherently challenging due to the intricate and often entangled relationships between various semantic features present within the images. However, we draw inspiration from a crucial aspect of our generative model: the latent space. This latent space, where latent vectors encode critical information about the data, offers a unique and powerful capability for rich and meaningful representation. It is within this latent space that we aim to excavate a robust representation of our undesired features. It is important to emphasize that the quality and effectiveness of this representation fundamentally depend on the inherent characteristics of the latent space itself. Consequently, the properties, dimensions, and organization of the latent space play a pivotal role in determining the accuracy and utility of the representation we derive.
@@
  $T_{th}\leftarrow\frac{\sum_{i=1}^{s}\text{sim}(z^{n}_{i},z_{undesired})+\text{sim}(z^{p}_{i},z_{undesired})}{2s}$
@@
IV-B2 Step-2 Similarity Thresholding
In the context of our work, we draw inspiration from the idea that latent vectors, which encode similar features, tend to be close to each other in the latent space of the generative model. This observation forms the basis for our approach, as we aim to effectively distinguish between latent vectors that correspond to desired and undesired features, ultimately allowing us to filter out the undesired outputs generated by the model. In this step, our approach involves the generation of a set of new latent vectors denoted as $\{z_{i}\}_{i=1}^{t}$. These latent vectors are sampled from a latent space distribution, typically following a normal distribution. Each $z_{i}$ is essentially a numerical representation of a potential data point that the generative model can produce. To assess the alignment between these newly sampled latent vectors and the undesired feature, we introduce a similarity metric, denoted as $\text{sim}(z,z_{undesired})$. This metric quantifies how similar a given latent vector $z$ is to the undesired feature vector $z_{undesired}$ in the latent space. The similarity score is computed as follows.
@@ IV-B2 Step-2 Similarity Thresholding
$T_{th}=\frac{\sum_{i=1}^{s}\text{sim}(z^{n}_{i},z_{undesired})+\text{sim}(z^{p}_{i},z_{undesired})}{2s}$ (15)
@@
Figure 2: FAST Filtering Mechanism: In Step-1, the undesired latent feature is identified using only a few positive and negative samples marked by the user. The positive and negative samples are projected into the latent space using the Latent Projection Function ($\pi(.)$), and subsequently, the undesired feature is retrieved via the Undesired Representation Function ($g(.)$). In Step-2, during the inference phase, new test samples are projected into the latent space, and the similarity of their projection with the undesired feature obtained from Step-1 is measured to filter out the negative samples.
IV-C Choices of LPF and URF
A crucial aspect to consider in our methodology is the flexibility and adaptability of the Latent Projection Function (LPF) and the Undesired Representation Function (URF). The Latent Projection Function (LPF) plays a vital role in mapping the data from its original space into the latent space, where different semantic features can be effectively encoded. On the other hand, the Undesired Representation Function (URF) is responsible for crafting a representation of the undesired feature within the latent space. These functions hold a pivotal role in the overall effectiveness of our approach, and selecting them appropriately is essential to ensure the method’s success. It’s worth emphasizing that there isn’t a one-size-fits-all solution; rather, the choice of these functions should be made based on the specific characteristics and demands of the problem under consideration.
Here we have analyzed three different LPFs that map into three different latent spaces as follows:
@@ IV-C Choices of LPF and URF
V Experiments and Results
V-A Datasets and Models
@@
  • Class-level Filtering: For this setting, we utilize the MNIST dataset (LeCun et al., 1998), which comprises 60,000 28 × 28 black and white images of handwritten digits. Aiming to achieve class-level filtering, we have used pre-trained DC-GAN on MNIST. Specifically, we focus on filtering out two digit classes: 5 and 8.
@@
  • Feature-level Filtering: In this scenario, we turn to the CelebA-HQ dataset (Liu et al., 2015), which contains 30,000 high-quality RGB celebrity face images with dimensions of 256×256. For this, we have taken state-of-the-art StyleGAN2 pre-trained of CelebA-HQ. Here, we target the feature-level filtering of subtle features, specifically (a) Bangs and (b) Hats.
V-B Training Details:
V-B1 Pre-trained GAN:
@@
As specified, we utilize two pre-trained GAN architectures: DCGAN trained on MNIST and Style-GAN2 trained on CelebA-HQ. The DCGAN architecture and weights are obtained from an open-source repository  [^1] https://github.com/csinva/gan-vae-pretrained-pytorch, while Style-GAN2 architecture is sourced from another repository  [^2] https://github.com/rosinality/stylegan2-pytorch. To obtain the pre-trained GAN on CelebA-HQ, images are resized to $256\times 256$ to conform to the StyleGAN2 architecture. The latent space dimensions are set to $100\times 1$ for MNIST and $512\times 1$ for CelebA-HQ. GAN training employs non-saturating adversarial loss with path-regularization. Default optimizers and hyperparameters provided in the code are utilized for training. The training duration for CelebA-HQ is $3.6\times 10^{5}$ epochs.
V-B2 Pretrained Classifier:
We leverage pre-trained classifiers to emulate user feedback, obtaining positive and negative samples by evaluating generated samples from the pre-trained GAN through these classifiers. Additionally, these classifiers simulate human filtering, and in the evaluation metrics section, we compute FID, density, and coverage based on this classifier’s output.
MNIST: We use simple LeNet model [44] for classification among different digits of MNIST dataset [^3] https://github.com/csinva/gan-vae-pretrained-pytorch/tree/master/mnist_classifier. The model is trained with a batch-size of $256$ using Adam optimizer with a learning rate of $2\times 10^{-3}$, $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The model is trained for a resolution of $32\times 32$ same as the pre-trained GAN for $12$ epochs. After training the classifier has an accuracy of $99.07\%$ on the test split of the MNIST dataset.
@@ V-B2 Pretrained Classifier:
CelebA-HQ: We use ResNext50 model [45] for classification among different facial attributes contained in CelebA [^4] https://github.com/rgkannan676/Recognition-and-Classification-of-Facial-Attributes/. Note that we train the classifier on normal CelebA as the ground truth values are available for it. The classifier is trained with a batch-size of $64$ using Adamax optimizer with a learning rate of $2\times 10^{-3}$, $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The model is trained for a resolution of $256\times 256$ for $10$ epochs. We also employ image augmentation techniques such as horizontal flip, image resize, and cropping to improve the performance of the classifier. The trained model exhibits a test accuracy of $91.93\%$.
V-B3 Base Classifier:
We employ a Base classifier trained on user feedback, utilizing a ResNext50 model [45] pretrained on Imagenet as the base feature extractor. The model consists of a 3-layer MLP trained with Binary Cross Entropy loss, fine-tuned for both MNIST and CelebA-HQ datasets. Adamax optimizer is used for training, running for 200 epochs with a learning rate of 0.07. Given the limited user feedback ($\leq 100$ samples), we introduce data augmentation techniques including rotation, blur, perspective, and auto-augment. Each image is augmented into 64 variations during training. Further details are provided in the Appendix section.
V-B4 Filtering Mechanism:
For the proposed filtering mechanism (Algorithm 1), we experimented with feature augmentation. Formally, we augment the number of negative and positive features in the latent space by sampling 5000 features from their respective empirical distributions, as given below:
Here, $\bar{z}^{p}$, $\bar{\Sigma}_{p}$, $\bar{z}^{n}$, $\bar{\Sigma}_{n})$ represent the mean and covariance matrix of the positive and negative latent vectors, respectively.
@@ V-B4 Filtering Mechanism:
TABLE I: Accuracy ($\uparrow$), Recall($\uparrow$), AUC($\uparrow$), FID ($\downarrow$) Density($\uparrow$) and Coverage($\uparrow$) after filtering MNIST classes with only 20 pos and 20 neg. user-feedback
Class-5 Class-8
Methods Acc. Rec. AUC FID Dens. Cov. Acc. Rec. AUC FID Dens. Cov.
Base Classifier 0.64$\pm$0.01 0.19$\pm$0.01 0.45$\pm$0.01 3.60$\pm$0.14 0.93$\pm$0.01 0.91$\pm$0.01 0.70$\pm$0.01 0.35$\pm$0.02 0.53$\pm$0.01 1.42$\pm$0.07 0.99$\pm$0.01 0.99$\pm$0.00
Base Classifier + Data Aug. 0.89$\pm$0.00 0.10$\pm$0.00 0.50$\pm$0.00 0.80$\pm$0.03 0.93$\pm$0.00 0.99$\pm$0.00 0.94$\pm$0.00 0.10$\pm$0.00 0.50$\pm$0.00 0.55$\pm$0.07 0.99$\pm$0.00 1.01$\pm$0.00
Imp-LS + MD 0.60$\pm$0.10 0.69$\pm$0.06 0.69$\pm$0.04 2.34$\pm$0.97 0.95$\pm$0.02 0.96$\pm$0.02 0.63$\pm$0.05 0.71$\pm$0.03 0.72$\pm$0.03 1.78$\pm$0.33 0.98$\pm$0.01 0.97$\pm$0.01
Imp-LS + MD + Latent Aug. 0.61$\pm$0.09 0.69$\pm$0.06 0.69$\pm$0.04 2.30$\pm$0.94 0.96$\pm$0.02 0.96$\pm$0.01 0.63$\pm$0.05 0.71$\pm$0.03 0.73$\pm$0.03 1.79$\pm$0.36 0.99$\pm$0.01 0.97$\pm$0.01
Imp-LS + Norm-SVM 0.61$\pm$0.11 0.67$\pm$0.11 0.69$\pm$0.03 2.27$\pm$1.00 0.96$\pm$0.01 0.96$\pm$0.02 0.62$\pm$0.05 0.69$\pm$0.04 0.71$\pm$0.02 1.90$\pm$0.36 0.99$\pm$0.02 0.97$\pm$0.01
Imp-LS + Norm-SVM + Latent Aug. 0.54$\pm$0.06 0.56$\pm$0.02 0.57$\pm$0.05 3.05$\pm$0.72 0.92$\pm$0.01 0.95$\pm$0.01 0.58$\pm$0.04 0.60$\pm$0.10 0.63$\pm$0.06 2.39$\pm$0.37 0.99$\pm$0.01 0.96$\pm$0.01
Inv-LS + MD 0.87$\pm$0.10 0.11$\pm$0.02 0.74$\pm$0.04 0.65$\pm$0.95 0.93$\pm$0.02 0.99$\pm$0.02 0.92$\pm$0.01 0.01$\pm$0.01 0.61$\pm$0.08 0.64$\pm$0.08 0.99$\pm$0.01 1.10$\pm$0.00
Inv-LS + MD + Latent Aug. 0.87$\pm$0.01 0.12$\pm$0.02 0.75$\pm$0.03 0.63$\pm$0.97 0.93$\pm$0.02 0.99$\pm$0.01 0.92$\pm$0..02 0.01$\pm$0.00 0.59$\pm$0.06 0.58$\pm$0.08 0.99$\pm$0.01 0.99$\pm$0.01
Inv-LS + Norm-SVM 0.85$\pm$0.13 0.55$\pm$0.02 0.83$\pm$0.04 0.72$\pm$1.01 0.96$\pm$0.01 0.97$\pm$0.02 0.89$\pm$0.06 0.11$\pm$0.14 0.68$\pm$0.04 0.85$\pm$0.33 0.99$\pm$0.00 0.99$\pm$0.01
Inv-LS + Norm-SVM + Latent Aug. 0.79$\pm$0.14 0.30$\pm$0.02 0.70$\pm$0.04 2.22$\pm$1.02 0.93$\pm$0.01 0.94$\pm$0.02 0.92$\pm$0.05 0.12$\pm$0.14 0.72$\pm$0.05 0.48$\pm$0.03 0.99$\pm$0.01 0.99$\pm$0.01
@@
TABLE II: Accuracy ($\uparrow$), Recall($\uparrow$), AUC($\uparrow$), FID ($\downarrow$) Density($\uparrow$) and Coverage($\uparrow$) after filtering Celeba-HQ features with only 20 pos and 20 neg. user-feedback

### v2

@@ 3.1 Machine Unlearning
(Exact Weak Unlearning) Given a retrained generative model $G_{\theta^{r}}(.)$, we say the model $G_{\theta^{u}}(.)$ is an exact weak unlearned model iff $\forall z\sim P_{\mathcal{Z}},\mathcal{O}\subset\mathcal{X}$ the following holds:
@@
(($\epsilon,\delta$) Approximate Weak Unlearning) Given a retrained generative model $G_{\theta^{r}}(.)$, we say the model $G_{\theta^{u}}(.)$ is an $(\epsilon,\delta)$ approximate weak unlearned model for a given $\epsilon,\delta\geq 0$ iff $\forall z\sim P_{\mathcal{Z}},\mathcal{O}\subset\mathcal{X}$ the following conditions hold:
@@
3.2 Filtering as Weak Unlearning
@@
Figure 1: Filtering as Weak Unlearning in Black-Box Generative Models: The left-most block represents the black-box generator with posthoc blocking layer with parameters $\theta^{b}=(\theta^{init},t)$ whose output distribution $P_{\theta^{b}}$, the middle block represents a retrained generative model with output distribution $P_{\theta^{r}}$, the right-most block represents a $(\epsilon,0)$ unlearned model with output distribution $P_{\theta^{u}}$.
@@
Remarks: Here the $d_{ln-TV}(.||.)$ denotes the total variational distance in terms of the log of the distributions.
Proof 3.2.
@@
Now, using ineq. (ii) and ineq. (iii) we get the below inequalities respectively
@@
4 Methodology
4.1 Problem Formulation and Method Overview
@@
As previously mentioned, the pre-trained generator $G_{\theta^{init}}$ is trained using a dataset $D$ comprising $m$ independent and identically distributed (i.i.d) samples denoted as $\{x_{i}\}_{i=1}^{m}$, where $x_{i}\overset{iid}{\sim}P_{\mathcal{X}}$. In our specific scenario, both the dataset $D$ and the initial model parameters $\theta^{init}$ remain undisclosed. Our objective is to formulate a filtering mechanism that relies on user feedback to selectively exclude undesired outputs. In this context, the user is presented with a collection of $r$ samples, denoted as $\mathcal{S}=\{y_{i}\}_{i=1}^{r}$, where each $y_{i}$ represents a generated sample originating from the pre-trained GAN. Within this user feedback system, the user’s role is to designate a relatively small subset of these samples, denoted as $\mathcal{S}_{n}=\{y^{n}_{i}\}_{i=1}^{s}$, as negative samples. These negative samples correspond to the instances that exhibit undesired features. Additionally, the user selects a portion of the samples, referred to as $\mathcal{S}_{p}=\{y^{p}_{i}\}_{i=1}^{s}$, as positive samples. Positive samples are those that do not possess the undesired features and are considered desirable. It’s essential to note that gathering user feedback can be resource-intensive. Therefore, we operate within a low-data regime, where the no of feedback samples marked by the user, denoted as $s$, is very small (constrained to be less than 100) i.e. $s\ll r$. This limitation acknowledges the potential challenges associated with obtaining extensive user feedback and emphasizes the need to devise effective filtering mechanisms even in scenarios with limited data availability.
@@
In this study, we implement a two-stage process to filter out undesired outputs effectively. In the initial stage termed as Undesired Feature Representation stage, we encode the representation of the undesired feature within the latent space. This latent space may either be the implicit latent space inherent to the Generative Adversarial Network (GAN) or a latent space that is learnable. Subsequently, in the second step termed as Similarity Thresholding stage, we compute the projection similarity between newly sampled latent vectors and this encoded representation of the undesired feature. This similarity calculation serves as the basis for determining the degree of alignment between the latent vectors and the undesired feature. Utilizing this similarity metric, we judiciously apply a thresholding mechanism to separate and exclude outputs that contain the undesired feature.
4.2 Filtering Mechanism
4.2.1 Undesired Feature Representation stage
@@
In this particular phase of our methodology, the central objective is to uncover a comprehensive and distinct representation of the undesired feature. This representation should exhibit a clear presence in the negative images while being absent in the positive images. This task is inherently challenging due to the intricate and often entangled relationships between various semantic features present within the images. However, as previously studied [35, 36, 38] the latent space, where latent vectors encode critical information about the data, offers a unique and powerful capability for rich and meaningful representation. But it is important to emphasize that the quality and effectiveness of this representation fundamentally depend on the inherent characteristics (e.g. disentanglement, clustering) of the latent space itself. Consequently, the properties, dimensions, and organization of the latent space play a pivotal role in determining the accuracy and utility of the representation we derive.
@@
  $T_{th}\leftarrow\frac{1}{2s}\sum_{i=1}^{s}\left[\text{sim}(z^{n}_{i},z_{undesired})+\text{sim}(z^{p}_{i},z_{undesired})\right]$
@@
4.2.2 Similarity Thresholding stage
In the context of our work, we draw inspiration from the clustering properties [38] of latent vectors, which encode similar features, tend to be close to each other in the latent space of the generative model. This observation forms the basis for our approach, as we aim to effectively distinguish between latent vectors that correspond to desired and undesired features, ultimately allowing us to filter out the undesired outputs generated by the model. In this step, our approach involves the generation of a set of new latent vectors denoted as $\{z_{i}\}_{i=1}^{t}$. These latent vectors are sampled from a latent space distribution, typically following a normal distribution. Each $z_{i}$ is essentially a numerical representation of a potential data point that the generative model can produce. To assess the alignment between these newly sampled latent vectors and the undesired feature, we introduce a similarity metric, denoted as $\text{sim}(z,z_{undesired})$. This metric quantifies how similar a given latent vector $z$ is to the undesired feature vector $z_{undesired}$ in the latent space. The similarity score is computed as follows.
@@ 4.2.2 Similarity Thresholding stage
$T_{th}=\frac{1}{2s}\sum_{i=1}^{s}\left[\text{sim}(z^{n}_{i},z_{undesired})+\text{sim}(z^{p}_{i},z_{undesired})\right]$ (15)
@@
Figure 2: FAST filtering mechanism: In stage-1, the undesired latent feature is identified using only a few positive and negative samples marked by the user. The positive and negative samples are projected into the latent space using the Latent Projection Function ($\pi(.)$), and subsequently, the undesired feature is retrieved via the Undesired Representation Function ($g(.)$). In stage-2, during the inference phase, new test samples are projected into the latent space, and the similarity of their projection with the undesired feature obtained from stage-1 is measured to filter out the negative samples.
4.3 Choices of LPF and URF
A crucial aspect to consider in our methodology is the flexibility and adaptability of the Latent Projection Function (LPF) and the Undesired Representation Function (URF). LPF plays a vital role in mapping the data from its original space into the latent space, where different semantic features can be effectively encoded. On the other hand, URF is responsible for crafting a representation of the undesired feature within the latent space. These functions hold a pivotal role in the overall effectiveness of our approach, and selecting them appropriately is essential to ensure the method’s success. It’s worth emphasizing that there isn’t a one-size-fits-all solution; rather, the choice of these functions should be made based on the specific characteristics and demands of the problem under consideration, making FAST a unified family of methods.
Here we have analyzed two different LPFs that map into two different latent spaces as follows:
@@ 4.3 Choices of LPF and URF
4.4 Latent Augmentation
The effectiveness of the proposed filtering mechanism is dependent on the limited feedback samples provided by the user. Due to the scarcity of user feedback samples, this method is prone to overfitting to the few available samples. To mitigate the overfitting issue and enhance the effectiveness of our method, we incorporate latent sampling in the latent space. Formally, we augment the number of negative and positive features in the latent space by sampling 5000 features from their respective empirical distributions, as given below:
Here, $\bar{z}^{p}$, $\bar{\Sigma}_{p}$, $\bar{z}^{n}$, $\bar{\Sigma}_{n}$ represent the mean and covariance matrix of the positive and negative latent vectors, respectively.
5 Experiments and Results
5.1 Datasets and Models
@@
  • Class-level filtering: For this setting, we utilize the MNIST dataset (LeCun et al., 1998), which comprises 60,000 black and white images of handwritten digits having dimensions of 28 × 28. Aiming to achieve class-level filtering, we have used pre-trained DC-GAN on MNIST. Specifically, we focus on filtering out two digit classes: 5 and 8.
@@
  • Feature-level filtering: In this scenario, we turn to the CelebA-HQ dataset (Liu et al., 2015), which contains 30,000 high-quality RGB celebrity face images with dimensions of 256×256. For this, we have taken state-of-the-art Style-GAN2 pre-trained on CelebA-HQ. Here, we target the feature-level filtering of subtle features, specifically (a) bangs and (b) hats.
5.2 Training Details:
5.2.1 Pre-trained GAN
@@
As specified, we utilize two pre-trained GAN architectures: DC-GAN trained on MNIST and Style-GAN2 trained on CelebA-HQ. The DC-GAN architecture and weights are obtained from an open-source repository  [^1] Source Code: Pre-trained DC-GAN trained on MNIST dataset, while Style-GAN2 architecture is sourced from another repository  [^2] Source Code: Pre-trained Style-GAN2 trained on CelebA-HQ dataset. To obtain the pre-trained GAN on CelebA-HQ, images are resized to $256\times 256$ to conform to the Style-GAN2 architecture. The latent space dimensions are set to $100\times 1$ for DC-GAN and $512\times 1$ for Style-GAN2. GAN training employs non-saturating adversarial loss with path-regularization. Default optimizers and hyperparameters provided in the code are utilized for training. The training duration for CelebA-HQ is $3.6\times 10^{5}$ epochs.
5.2.2 Pre-trained Classifier
@@
We leverage pre-trained classifiers to emulate user feedback, obtaining positive and negative samples by evaluating generated samples from the pre-trained GAN through these classifiers. Additionally, these classifiers simulate human filtering, and in the evaluation metrics section, we compute FID, density, and coverage based on this classifier’s output. We use simple LeNet model [^3] Source Code: Pre-trained Classifier trained on MNIST dataset for classification among different digits of MNIST dataset. The model is trained with a batch-size of $256$ using Adam optimizer with a learning rate of $2\times 10^{-3}$, $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The model is trained for a resolution of $32\times 32$ same as the pre-trained GAN for $12$ epochs. After training the classifier has an accuracy of $99.07\%$ on the test split of the MNIST dataset. Further for feature level-filtering, we use ResNext50 model [44] for classification among different facial attributes contained in CelebA-HQ [^4] Source Code: Pre-trained Classifier trained on CelebA-HQ dataset. The classifier is trained with a batch-size of $64$ using Adamax optimizer with a learning rate of $2\times 10^{-3}$, $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The model is trained for a resolution of $256\times 256$ for $10$ epochs. We also employ image augmentation techniques such as horizontal flip, image resize, and cropping to improve the performance of the classifier. The trained model exhibits a test accuracy of $91.93\%$.
@@
Table 1: Recall($\uparrow$), AUC($\uparrow$), FID ($\downarrow$) Density($\uparrow$) and Coverage($\uparrow$) after filtering MNIST classes with only 20 Pos. and 20 Neg. user-feedback
Methods Class-5 Class-8
Recall AUC FID Density Coverrage Recall AUC FID Density Coverage
Base Classifier 0.19$\pm$0.01 0.45$\pm$0.01 3.60$\pm$0.14 0.93$\pm$0.01 0.91$\pm$0.01 0.35$\pm$0.02 0.53$\pm$0.01 1.42$\pm$0.07 0.99$\pm$0.01 0.99$\pm$0.00
Base Classifier+Data Aug. 0.10$\pm$0.00 0.50$\pm$0.00 0.80$\pm$0.03 0.93$\pm$0.00 0.99$\pm$0.00 0.10$\pm$0.00 0.50$\pm$0.00 0.55$\pm$0.07 0.99$\pm$0.00 1.01$\pm$0.00
Imp-LS+MD 0.69$\pm$0.06 0.69$\pm$0.04 2.34$\pm$0.97 0.95$\pm$0.02 0.96$\pm$0.02 0.71$\pm$0.03 0.72$\pm$0.03 1.78$\pm$0.33 0.98$\pm$0.01 0.97$\pm$0.01
Imp-LS+MD+Latent Aug. 0.69$\pm$0.06 0.69$\pm$0.04 2.30$\pm$0.94 0.96$\pm$0.02 0.96$\pm$0.01 0.71$\pm$0.03 0.73$\pm$0.03 1.79$\pm$0.36 0.99$\pm$0.01 0.97$\pm$0.01
Imp-LS+Norm-SVM 0.67$\pm$0.11 0.69$\pm$0.03 2.27$\pm$1.00 0.96$\pm$0.01 0.96$\pm$0.02 0.69$\pm$0.04 0.71$\pm$0.02 1.90$\pm$0.36 0.99$\pm$0.02 0.97$\pm$0.01
Imp-LS+Norm-SVM+Latent Aug. 0.56$\pm$0.02 0.57$\pm$0.05 3.05$\pm$0.72 0.92$\pm$0.01 0.95$\pm$0.01 0.60$\pm$0.10 0.63$\pm$0.06 2.39$\pm$0.37 0.99$\pm$0.01 0.96$\pm$0.01
Inv-LS+MD 0.11$\pm$0.02 0.74$\pm$0.04 0.65$\pm$0.95 0.93$\pm$0.02 0.99$\pm$0.02 0.01$\pm$0.01 0.61$\pm$0.08 0.64$\pm$0.08 0.99$\pm$0.01 1.10$\pm$0.00
Inv-LS+MD+Latent Aug. 0.12$\pm$0.02 0.75$\pm$0.03 0.63$\pm$0.97 0.93$\pm$0.02 0.99$\pm$0.01 0.01$\pm$0.00 0.59$\pm$0.06 0.58$\pm$0.08 0.99$\pm$0.01 0.99$\pm$0.01
Inv-LS+Norm-SVM 0.55$\pm$0.02 0.83$\pm$0.04 0.72$\pm$1.01 0.96$\pm$0.01 0.97$\pm$0.02 0.11$\pm$0.14 0.68$\pm$0.04 0.85$\pm$0.33 0.99$\pm$0.00 0.99$\pm$0.01
Inv-LS+Norm-SVM+Latent Aug. 0.30$\pm$0.02 0.70$\pm$0.04 2.22$\pm$1.02 0.93$\pm$0.01 0.94$\pm$0.02 0.12$\pm$0.14 0.72$\pm$0.05 0.48$\pm$0.03 0.99$\pm$0.01 0.99$\pm$0.01
@@
Table 2: Recall($\uparrow$), AUC($\uparrow$), FID ($\downarrow$) Density($\uparrow$) and Coverage($\uparrow$) after filtering Celeba-HQ features with only 20 Pos. and 20 Neg. user-feedback